# CuckooGraph: A Scalable and Space-Time Efficient Data Structure for Large-Scale Dynamic Graphs

### v2

@@ I-B Our Proposed Solution
In this paper, we propose a novel data structure for storing large-scale dynamic graphs, namely CuckooGraph [^1] Note that CuckooGraph is essentially an algorithm-level solution that pursues faster insertion and query speed and less memory usage to serve graph-related tasks, different from those system-level solutions.. It has the following advantages: 1) It is memory-saving and can be flexibly expanded or contracted according to actual operations; 2) It basically maintains the fastest running speed in a series of graph analytics tasks; 3) It works for any graph of unknown size without knowing any information about it in advance.
@@
The rest of this paper is organized as follows. We theoretically prove that the time and memory cost of CuckooGraph is desirable through mathematical analysis in §$\S$ IV. We conduct extensive experiments on 7 large-scale graph datasets with different characteristics to evaluate the performance of CuckooGraph on basic tasks and graph analytics tasks. The results clearly show that CuckooGraph has the fastest speed and the lowest memory overhead on almost all tasks, see §$\S$ V for details.
@@ II Related Work
In this section, we introduce graph storage schemes whose main contribution lies in data structure design and Cuckoo Hash Table (CHT)—the most basic data structure of CuckooGraph, in §$\S$ II-A and §$\S$ II-B, respectively.
@@ II-A Existing Solutions
II-B Cuckoo Hashing
@@ III-A Basic Version
III-A1 Transformable Data Structures
As shown in Figure 1 (except for the rightmost orange box), the basic data structure of CuckooGraph consists of one (or more) large cuckoo hash table(s) (denoted as L-CHT(s)) and many small cuckoo hash tables (denoted as S-CHTs) associated with it/them, all of which are specially designed.
L-CHT has two bucket arrays denoted as $B_{1}$ and $B_{2}$, respectively, associated with two independent hash functions $H_{1}(.)$ and $H_{2}(.)$, respectively [^3] The basic structure of S-CHT is no different from that of L-CHT, except for the length and the association with two other independent hash functions $h_{1}(.)$ and $h_{2}(.)$.. Each bucket has $d$ cells, each of which is designed to have two parts: Part 1 and Part 2. For any arriving graph item $\left\langle u,v\right\rangle$, assuming it is mapped to bucket $B_{H}$ and deposited in the cell $C$, then: Part 1 of $C$ is used to store $u$, while Part 2 of $C$ is directly used to store $v$ or the pointer(s) pointing to S-CHT which is used to actually store $v$. In Figure 1 (red box), Part 2 is designed as a structure that can be flexibly transformed in a manner determined by the number of $v$ (denoted as $l$) corresponding to $u$ in Part 1, as shown below: ① Part 2 is initialized to $2R$ small slots, $i.e.$, up to $2R$ $v$ can be recorded, to handle the situation where $l\leq 2R$; ② When $l>2R$, $2R$ small slots are merged in pairs to form $R$ large slots dedicated to storing $R$ pointers usually with more bytes, and 1st large slot is deposited with a pointer that points to 1st S-CHT; Then, all the current $v$ are stored into this S-CHT of length [^4] We define the length of CHT as the number of buckets in the array with more buckets. $n$.
@@ III-A1 Transformable Data Structures
III-A2 Denylist Optimization
There is one aspect of our design that has not been considered so far: the original CHT may suffer from item insertion failures. A straightforward solution is to extend CuckooGraph via the transformations described above whenever an insertion failure occurs. To address this issue more efficiently, we further propose an optimization called Denylist (DL), as shown in Figure 1 (orange box).
@@ IV-A Time Cost of CuckooGraph
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}2Gn+4Gn+8Gn+\cdots+2^{k}Gn=2(2^{k}-1)Gn},$
@@ V Evaluation
In this section, we evaluate the performance of CuckooGraph through extensive experiments, which are briefly described as follows: 1) We introduce the experimental setup in §$\S$ V-A; 2) We evaluate how key parameters affect CuckooGraph in §$\S$ V-B; 3) We verify the effect of Denylist optimization through ablation experiments in §$\S$ V-C; 4) We evaluate the insertion, query, and deletion throughput as well as memory usage of CuckooGraph and its competitors in §$\S$ V-D; 5) We evaluate the running time of CuckooGraph and its competitors on graph analytics tasks (BFS, SSSP, TC, CC, PR, BC, LCC) in §$\S$ V-E.
TABLE III: A brief analysis of the graph datasets used.
@@ V-A Experimental Setup
Datasets: We use various graph datasets to comprehensively evaluate the performance of CuckooGraph and its competitors, and the details are shown in Table III. 1) The CAIDA dataset [46] is streams of anonymized IP traces collected by CAIDA. Each flow is identified by a five-tuple: source and destination IP addresses, source and destination ports, protocol. The source and destination IP addresses in the traces are used as the start and end nodes of the graphs, respectively. 2) The NotreDame dataset [47] is a web graph collected from University of Notre Dame. Nodes represent web pages, and directed edges represent hyperlinks between them. 3) The StackOverflow dataset [48] is a collection of interactions on the stack exchange website called Stack Overflow. Nodes represent users and edges represent user interactions. 4) The WikiTalk dataset [49] is a collection of user communications obtained from English Wikipedia [49], and the nodes and edges refer to the same as above. 5) The Weibo dataset [50] is captured from Sina Weibo Open Platform APIs, and the definitions of nodes and edges are similar to those of StackOverflow. 6) We synthesize the DenseGraph dataset. 7) We synthesize the SparseGraph dataset.
@@ V-E Experiments on Graph Analytics Tasks
In this subsection, we evaluate the performance of CuckooGraph and its competitors in terms of running time on the graph datasets in Table III through the following typical graph analytics tasks: Breadth-First Search (BFS), Single-Source Shortest Paths (SSSP), Triangle Counting (TC), Connected Components (CC), PageRank (PR), Betweenness Centrality (BC), and Local Clustering Coefficient (LCC). Note that some competitors did not complete the experiments within the given time, so their results are not shown in the provided figures.
@@ V-E1 Breadth-First Search
Figure 10: Running time of BFS on different datasets.
Figure 11: Running time of SSSP on different datasets.
Figure 12: Running time of TC on different datasets.
Figure 13: Running time of CC on different datasets.
@@ V-E3 Triangle Counting
Figure 14: Running time of PR on different datasets.
Figure 15: Running time of BC on different datasets.
Figure 16: Running time of LCC on different datasets.
@@ VI Conclusion
In this paper, we propose a novel data structure designed for large-scale dynamic graphs, called CuckooGraph, which includes two key techniques, Transformation and Denylist. Thanks to them, CuckooGraph can be flexibly resized based on actual operations to achieve memory efficiency while keeping few memory accesses to achieve fast processing speed without any prior knowledge of the upcoming graphs. Our mathematical analysis theoretically proves that CuckooGraph is time and space efficient. Our experimental results show that CuckooGraph significantly outperforms 4 state-of-the-art schemes. In particular, compared with Spruce, CuckooGraph achieves 32.66×32.$6$ faster insertion throughput while reducing memory space by about $32\%$, and 168.45×168.45$\times$168.45 × less running time on the SSSP task.

### v3

@@ I-B Our Proposed Solution
In this paper, we propose a novel data structure for storing large-scale dynamic graphs, namely CuckooGraph. It has the following advantages: 1) It is memory-saving and can be flexibly expanded or contracted according to actual operations; 2) It basically maintains the fastest running speed in a series of graph analytics tasks; 3) It works for any graph of unknown size without knowing any information about it in advance.
@@
The rest of this paper is organized as follows. We theoretically prove that the time and memory cost of CuckooGraph is desirable through mathematical analysis in §$\S$ IV. We conduct extensive experiments on 7 large-scale graph datasets with different characteristics to evaluate the performance of CuckooGraph on basic tasks and graph analytics tasks. The results clearly show that CuckooGraph has the fastest speed and the lowest memory overhead on almost all tasks. Finally, CuckooGraph is integrated into Redis and Neo4j databases as use cases. See §$\S$ V for more details. All relevant source code is already open source on Github [43].
@@ II Related Work
In this section, we introduce graph storage schemes whose main contribution lies in data structure design and Cuckoo Hash Table (CHT)—the most basic data structure of CuckooGraph, in §$\S$ II-A and §$\S$ II-C, respectively.
@@ II-A Existing Solutions
II-B Orthogonal Work
Recently, there are notable works such as VEND [46] that aim to accelerate edge queries by filtering no-result edges, which are orthogonal to graph storage/databases. Since most node pairs in the real world have only no-edge connections, VEND introduces a novel data structure for nodes to store redundant neighbor information based on range and hash solutions. We leave the possibility of applying VEND to CuckooGraph as future work.
II-C Cuckoo Hashing
@@ III-A Basic Version
III-A1 Foundation Stage (Transformable Data Structures)
As shown in Figure 1 (Foundation Stage), the basic data structure of CuckooGraph consists of one (or more) large cuckoo hash table(s) (denoted as L-CHT(s)) and many small cuckoo hash tables (denoted as S-CHTs) associated with it/them, all of which are specially designed.
L-CHT has two bucket arrays denoted as $B_{1}$ and $B_{2}$, respectively, associated with two independent hash functions $H_{1}(.)$ and $H_{2}(.)$, respectively [^2] The basic structure of S-CHT is no different from that of L-CHT, except for the length and the association with two other independent hash functions $h_{1}(.)$ and $h_{2}(.)$.. Each bucket has $d$ cells, each of which is designed to have two parts: Part 1 and Part 2. For any arriving graph item $\left\langle u,v\right\rangle$, assuming it is mapped to bucket $B_{H}$ and deposited in the cell $C$, then: Part 1 of $C$ is used to store $u$, while Part 2 of $C$ is directly used to store $v$ or the pointer(s) pointing to S-CHT which is used to actually store $v$. In Foundation Stage (red box), Part 2 is designed as a structure that can be flexibly transformed in a manner determined by the number of $v$ (denoted as $l$) corresponding to $u$ in Part 1, as shown below: ① Part 2 is initialized to $2R$ small slots, $i.e.$, up to $2R$ $v$ can be recorded, to handle the situation where $l\leq 2R$; ② When $l>2R$, $2R$ small slots are merged in pairs to form $R$ large slots dedicated to storing $R$ pointers usually with more bytes, and 1st large slot is deposited with a pointer that points to 1st S-CHT; Then, all the current $v$ are stored into this S-CHT of length [^3] We define the length of CHT as the number of buckets in the array with more buckets. $n$.
@@ III-A1 Foundation Stage (Transformable Data Structures)
III-A2 Optimization Stage (Denylist)
There is one aspect of our design that has not been considered so far: the original CHT may suffer from item insertion failures. A straightforward solution is to extend CuckooGraph via the transformations described above whenever an insertion failure occurs. To address this issue more efficiently, we further propose an optimization called Denylist (DL), as shown in Figure 1 (Optimization Stage).
@@ IV-A Time Cost of CuckooGraph
TABLE III: Comparison of complexity between different solutions.
Algorithm Amortized Time Complexity Space Complexity
Insert Edge $\langle u,v\rangle$ Query Edge $\langle u,v\rangle$
LiveGraph [30] $O(1)$ $O(\deg(v))$ $O(|E|)$
Spruce [36] $O\left(\frac{|E|}{|V|}\right)$ $O\left(\log\frac{|E|}{|V|}\right)$ $O(|E|)$
Sortledton [34] $O(\log|E|)$ $O(\log|E|)$ $O(|E|)$
WBI [35] $O(1)$ $O(\frac{|E|}{K^{2}})$ $O(K^{2}+|E|)$
CuckooGraph (Ours) $O(1)$ $O(1)$ $O(|E|)$
@@
$2Gn+4Gn+8Gn+\cdots+2^{k}Gn=2(2^{k}-1)Gn,$
@@ IV-B Memory Cost of CuckooGraph
IV-C Discussions
In this part, we provide Table III to summarize the time and space complexities of CuckooGraph and some state-of-the-art schemes ($i.e.$, competitors mentioned in §$\S$ V). Here, $K$ refers to the length/width of the matrix, which is a parameter of WBI. In summary, our CuckooGraph has an insertion and query time complexity of $O(1)$ when Theorem 1 holds and $T$ is not very large, while its space complexity is still $O(|E|)$. The previous analysis in §$\S$ IV-B also proves that its space overhead is very small.
@@ V Evaluation
In this section, we evaluate the performance of CuckooGraph through extensive experiments, which are briefly described as follows: 1) We introduce the experimental setup in §$\S$ V-A; 2) We evaluate how key parameters affect CuckooGraph in §$\S$ V-B; 3) We verify the effect of Denylist optimization through ablation experiments in §$\S$ V-C; 4) We evaluate the insertion, query, and deletion throughput as well as memory usage of CuckooGraph and its competitors in §$\S$ V-D; 5) We evaluate the running time of CuckooGraph and its competitors on graph analytics tasks (BFS, SSSP, TC, CC, PR, BC, LCC) in §$\S$ V-E; 6) We deploy CuckooGraph on Redis and Neo4j databases and evaluate the speed in §$\S$ V-F and §$\S$ V-G, respectively.
TABLE IV: A brief analysis of the graph datasets used.
@@ V-A Experimental Setup
Datasets: We use various graph datasets to comprehensively evaluate the performance of CuckooGraph and its competitors, and the details are shown in Table IV. 1) The CAIDA dataset [49] is streams of anonymized IP traces collected by CAIDA. Each flow is identified by a five-tuple: source and destination IP addresses, source and destination ports, protocol. The source and destination IP addresses in the traces are used as the start and end nodes of the graphs, respectively. 2) The NotreDame dataset [50] is a web graph collected from University of Notre Dame. Nodes represent web pages, and directed edges represent hyperlinks between them. 3) The StackOverflow dataset [51] is a collection of interactions on the stack exchange website called Stack Overflow. Nodes represent users and edges represent user interactions. 4) The WikiTalk dataset [52] is a collection of user communications obtained from English Wikipedia, and the nodes and edges refer to the same as above. 5) The Weibo dataset [53] is captured from Sina Weibo Open Platform APIs, and the definitions of nodes and edges are similar to those of StackOverflow. 6) We synthesize the DenseGraph dataset. 7) We synthesize the SparseGraph dataset.
@@ V-E Experiments on Graph Analytics Tasks
In this subsection, we evaluate the performance of CuckooGraph and its competitors in terms of running time on the graph datasets in Table IV through the following typical graph analytics tasks: Breadth-First Search (BFS), Single-Source Shortest Paths (SSSP), Triangle Counting (TC), Connected Components (CC), PageRank (PR), Betweenness Centrality (BC), and Local Clustering Coefficient (LCC). Note that some competitors did not complete the experiments within the given time, so their results are not shown in the provided figures.
@@ V-E2 Single-Source Shortest Paths
Figure 11: Running time of SSSP on different datasets.
@@ V-E7 Local Clustering Coefficient
V-F Redis Implementation
Methodology: We utilize Redis Module [58] to register our CuckooGraph module, adding the data structure of CuckooGraph to the original Redis. This allows Redis to store graphs in addition to supporting the five original data structures. Specifically, we implement Redis Module API (including save_rdb, load_rdb, aof_rewrite and other interfaces) on top of CuckooGraph to support Redis persistence operations. Meanwhile, we also provide extended commands for CuckooGraph (including insert, del, query and getneighbors). We compile our interface implementation into a dynamic link library, and simply import CuckooGraph library with --loadmodule when Redis starts.
Setup: We conduct experiments on the CAIDA and StackOverflow datasets to test the throughput performance of CuckooGraph on Redis.
Results & Analysis (Figure 17): The results show that the insertion, query, and deletion throughput of CuckooGraph on Redis is around $0.04\sim 0.05$ Mops. There is some performance loss compared to those of CuckooGraph on CPU, which is mainly caused by the Redis system. We also run Redis benchmark on the server, and the peak throughput of native Redis is only around $0.16$ Mops. Considering that CuckooGraph itself inevitably has overhead, its performance on Redis is completely acceptable.
Figure 17: The throughput of CuckooGraph on Redis.
V-G Neo4j Implementation
Methodology: If we want to store an edge $\left\langle u,v\right\rangle$ in Neo4j [37], nodes $u$ and $v$ each maintain a adjacency list that stores all the edges associated with that node, so the information about $\left\langle u,v\right\rangle$ is stored in the adjacency lists of both $u$ and $v$. If we want to query an edge $\left\langle u,v\right\rangle$, we have to find the adjacency list of $u$, and then traverse the list and compare the edges one by one until we find $\left\langle u,v\right\rangle$. Obviously, it is inefficient. Once the degree of $u$ is high, querying edge $\left\langle u,v\right\rangle$ has to access a large number of unrelated other edges, causing additional redundancy overhead. To speed up edge queries, we introduce the CuckooGraph query interface to obtain an edge without traversing the adjacency list of the node. Since multiple edges (with the same $u$ and $v$ but not the same edge) are allowed in Neo4j, the data structure of CuckooGraph needs some adjustments for this. Compared to the weighted version on the CPU, we change the weight field in each S-CHT small slot from a counter that records the number of edges to a linked list consisting of a series of edges with the same nodes $u$ and $v$. The linked list is as long as the number of edges corresponding to $\left\langle u,v\right\rangle$ in that small slot. In this way, the query interface of CuckooGraph returns an iterator, through which the linked list can be traversed to obtain all the edges between $\left\langle u,v\right\rangle$.
Setup: We deploy the above CuckooGraph on top of the original Neo4j and evaluate the performance by running time, as shown below. 1) For the insertion experiments, we insert the first 1M edges from the CAIDA dataset into Neo4j. Whenever an edge is inserted into Neo4j, we also need to insert that edge into the CuckooGraph structure, which requires a little extra time overhead. 2) For the query experiments, we first deduplicate the 1M edges, and then query the CuckooGraph structure. For comparison, related operations on pure Neo4j do not introduce CuckooGraph.
Results & Analysis (Figure 18): 1) Thanks to the good performance of our data structure, our insertions are very fast and require only a little extra overhead, so our insertion time is almost the same as pure Neo4j. 2) Since the time cost of CuckooGraph’s query to obtain the iterator of the linked list is $O(1)$, the query speed of the version with CuckooGraph is very fast. It can be predicted that the query speed of Neo4j with CuckooGraph will be improved more significantly as the data scale increases. In pure Neo4j, many irrelevant/redundant edges must be traversed, and this additional overhead time is not $O(1)$, which ultimately causes the query time of pure Neo4j to be much slower than that with the assistance of CuckooGraph.
Figure 18: Running time of Neo4j with and without CuckooGraph.
@@ VI Conclusion
In this paper, we propose a novel data structure designed for large-scale dynamic graphs, called CuckooGraph, which includes two key techniques, Transformation and Denylist. Thanks to them, CuckooGraph can be flexibly resized based on actual operations to achieve memory efficiency while keeping few memory accesses to achieve fast processing speed without any prior knowledge of the upcoming graphs. Our mathematical analysis theoretically proves that CuckooGraph is time and space efficient. Our experimental results show that CuckooGraph significantly outperforms 4 state-of-the-art schemes. In particular, compared with Spruce, CuckooGraph achieves 32.66×32.$6$ faster insertion throughput while reducing memory space by about $32\%$, and 168.45×168.45$\times$168.45 × less running time on the SSSP task. Finally, we integrate CuckooGraph in Redis and Neo4j databases to extend its practicality.
Acknowledgment
We would like to thank the anonymous reviewers for their insightful comments. This work was supported by the National Natural Science Foundation of China (NSFC) (No. 62402012), and in part by the China Postdoctoral Science Foundation (No. 2023TQ0010, GZC20230055, 2024M750102).